OpenStudy (tommy14):

i need help whats 5x4

OpenStudy (joker):

5x4 as in multiplication? Add five four time or four five times. 5 + 5 + 5 + 5 or 4 + 4 + 4 + 4 + 4

OpenStudy (mathstudent55):

Multiplication is repeated addition. 5 * 4 = 4 * 5 = 5 + 5 + 5 + 5 If you have difficult with that addition, use your fingers and toes.

OpenStudy (hayley):

20

OpenStudy (mathstudent55):

Correct.

OpenStudy (codyisamazing):

you really dont know what 5*4 is?